# Comparing Two Proxy Methods for Causal Identification

### v1

@@ 1 Introduction
The array decomposition approach imposes similar invertibility conditions, but rather than solving directly for a causal target, it recovers the full law of latent and observed variables, from which causal effects follow. Originating from the work of Kruskal (1977), this framework allows causal effect identification in models with arbitrary hidden variables, such as in settings with unmeasured confounders (Kuroki and Pearl, 2014) or unmeasured exposures (Zhou and Tchetgen, 2024). Though initially formulated for effect identification in fully discrete models (Kuroki and Pearl, 2014; Allman, Matias and Rhodes, 2009), its identifying assumptions extend to continuous settings via the results of Hu and Schennach (2008), as shown by Deaner (2023).
Despite their shared ability to identify causal effects in the presence of unmeasured variables, these two approaches exhibit non-overlapping model restrictions. Prior work from Deaner (2023) has shown that the conditional independence assumptions (i.e., Markov restrictions) underlying the two approaches for hidden confounding are distinct and the models non-nested.
@@ 2.0.2 Solution to a Bridge Equation
where $\mathbf{P}_{W\mid Z,a}^{\dagger}$ denotes a left inverse of $\mathbf{P}_{W\mid Z,a}$. This requires $W$ to have at least as many categories as $Z$, and therefore at least as many as the latent variable $U$ under Assumption 6.
@@ 3 Causal Target Identification With Proxies via Array Decomposition
Another prominent strategy for identifying counterfactual quantities with proxies of hidden variables employs array decomposition to recover the full law of hidden and observed variables, up to the labels of hidden variable categories (e.g., Kuroki and Pearl (2014); Allman et al. (2015); Zhou and Tchetgen (2024)). We now review the assumptions of this approach in the context of the discrete hidden confounding model depicted in Figure 3, from Kuroki and Pearl (2014). In Appendix B, we provide analogous assumptions which leverage proxies of a hidden mediator in the absence of hidden confounder proxies, enabling identification causal effects via the frontdoor adjustment formula in Equation 7.
@@
A detailed proof of identification is given in Kuroki and Pearl (2014). For Assumption 10 to be satisfied, $Z$ and $W$ must have the same number of levels. Assumption 10 can be relaxed to left invertibility of $\mathbf{P}_{Z\mid U,a}$ and $\mathbf{P}_{W\mid U}$, allowing cases where $Z$ and $W$ each have at least as many levels as $U$. Perhaps surprisingly, Assumption 10 can even be relaxed to allow settings where both proxies $Z$ and $W$ have fewer levels than the latent variable. These generalizations are discussed in Section 3.1.
@@ 4.0.1 Discrete Model Comparison
For discrete models, we compare Assumption 6 and 7 from the bridge equation approach in Miao et al. (2018) with two alternative assumption sets. Set 1 is from Kuroki and Pearl (2014); Set 2 is from Kruskal (1977), the generalization of Kuroki and Pearl (2014) via Theorem 3.5.
@@ 4.0.2 Continuous Model Comparison
For $g$ square-integrable, $\mathbb{E}[g(u)\mid z,a]=0\;\mathrm{a.s.}\iff g(u)=0\;\mathrm{a.s.}$
@@
For $g$ square-integrable, $\mathbb{E}[g(u)\mid z,a]=0\;\mathrm{a.s.}\iff g(u)=0\;\mathrm{a.s.}$ and $\mathbb{E}[g(u)\mid w,a]=0\;\mathrm{a.s.}\iff g(u)=0\;\mathrm{a.s.}$
@@ Appendix A Frontdoor Adjustment Formula
Appendix B Mediator Proxy Analogue for Kuroki and Pearl (2018)
@@
We prove $f_{Y(a)}(y)$ is identified, relying on Lemma B.1.
Lemma B.1.
$\mathbf{M}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{-1},$
Proof.
$\mathbf{P}_{W\mid Z,a}=\mathbf{P}_{W\mid M}\mathbf{P}_{M\mid Z,a},\text{ and}$
$\mathbf{P}_{y,W\mid Z,a}=\mathbf{P}_{W\mid M}\operatorname{diag}(\mathbf{P}_{y\mid M,a})\mathbf{P}_{M\mid Z,a}.$
By Assumption 27, $\mathbf{P}_{M,Z\mid a}^{-1}$ exists and
$\mathbf{P}_{W\mid Z,a}^{-1}=\mathbf{P}_{M\mid Z,a}^{-1}\,\mathbf{P}_{W\mid M}^{-1}.$
Hence, $\mathbf{P}_{y,W\mid Z,a}\mathbf{P}_{W\mid Z,a}^{-1}=\mathbf{P}_{W\mid M}\operatorname{diag}(\mathbf{P}_{y\mid M,a})\mathbf{P}_{W\mid M}^{-1}.$
Lemma B.1 allows us to identify the columns of $\mathbf{P}_{W\mid M}$ under Assumptions 27 and 28. In other words, we recover conditional distributions $f_{W\mid M}(w\mid m_{i})$ up to labels $i$. Inverting the integral equation $f_{Y,W\mid A}(y,w\mid a)=\sum_{i}f_{Y,M\mid A}(y,m_{i}\mid a)\,f_{W\mid M}(w\mid m_{i}),$ we can solve for $f_{Y,M\mid A}(y,m_{i}\mid a)$ for each $a\in\mathcal{A}$ up to labeling. Further by $W\perp M\mid A$, we can align latent factors for label $i$ across values of $a\in\mathcal{A}$, enabling reconstruction of the frontdoor adjustment formula in Equation 7. ∎

### v2

@@ 1 Introduction
The array decomposition approach imposes similar invertibility conditions, but rather than solving directly for a causal target, it recovers the full law of latent and observed variables, from which causal effects follow. Originating from the work of Kruskal (1977), this framework allows causal effect identification in models with arbitrary hidden variables, such as in settings with unmeasured confounders (Kuroki and Pearl, 2014) or unmeasured exposures (Zhou and Tchetgen, 2024). Though initially formulated for effect identification in fully discrete models (Kuroki and Pearl, 2014; Allman et al., 2015), its identifying assumptions extend to continuous settings via the results of Hu and Schennach (2008), as shown by Deaner (2023).
Despite their shared ability to identify causal effects in the presence of unmeasured variables, these two approaches exhibit non-nested model restrictions. Prior work from Deaner (2023) has shown that the conditional independence assumptions (i.e., Markov restrictions) underlying the two approaches for hidden confounding are distinct and the models non-nested.
@@ 2.0.2 Solution to a Bridge Equation
where $\mathbf{P}_{W\mid Z,a}^{\dagger}$ denotes a left inverse of $\mathbf{P}_{W\mid Z,a}$. This requires $W$ to have at least as many categories as $Z$, and therefore at least as many as the latent variable $U$ under Assumption 6. From Assumption 6,
$\mathbf{P}_{Y\mid U,a}=\mathbf{P}_{Y\mid Z,a}\,\mathbf{P}_{W\mid Z,a}^{\dagger}\,\mathbf{P}_{W\mid U}.$
Then taking the inner product with $f_{U}(u)$ on both sides gives $f_{Y(a)}(y)$.
@@ 3 Causal Target Identification With Proxies via Array Decomposition
Another prominent strategy for identifying counterfactual quantities with proxies of hidden variables employs array decomposition to recover the full law of hidden and observed variables, up to the labels of hidden variable categories (e.g., Kuroki and Pearl (2014); Allman et al. (2015); Zhou and Tchetgen (2024)). We now review the assumptions of this approach in the context of the discrete hidden confounding model depicted in Figure 3, from Kuroki and Pearl (2014). In Appendix C, we provide analogous assumptions which leverage proxies of a hidden mediator in the absence of hidden confounder proxies, enabling identification causal effects via the frontdoor adjustment formula in Equation 7.
@@
A detailed proof of identification is given in Kuroki and Pearl (2014). We provide a short summary in Appendix B. For Assumption 10 to be satisfied, $Z$ and $W$ must have the same number of levels. Assumption 10 can be relaxed to left invertibility of $\mathbf{P}_{Z\mid U,a}$ and $\mathbf{P}_{W\mid U}$, allowing cases where $Z$ and $W$ each have at least as many levels as $U$. Perhaps surprisingly, Assumption 10 can even be relaxed to allow settings where both proxies $Z$ and $W$ have fewer levels than the latent variable. These generalizations are discussed in Section 3.1.
@@ 4.0.1 Discrete Model Comparison
For discrete models, we compare Assumption 6 and 7 from the bridge equation approach in Miao et al. (2018) with two alternative assumption sets (Table 1). Set 1 is from Kuroki and Pearl (2014); Set 2 is from Kruskal (1977), the generalization of Kuroki and Pearl (2014) via Theorem 3.5.
@@ 4.0.2 Continuous Model Comparison
[…]
$\mathbb{E}[g(u)\mid z,a]=0\;\mathrm{a.s.}\iff g(u)=0\;\mathrm{a.s.}$
@@
[…]
$\mathbb{E}[g(u)\mid z,a]=0\;\mathrm{a.s.}\iff g(u)=0\;\mathrm{a.s.}$ and $\mathbb{E}[g(u)\mid w,a]=0\;\mathrm{a.s.}\iff g(u)=0\;\mathrm{a.s.}$
@@ Appendix A Frontdoor Adjustment Formula
Appendix B Kuroki and Pearl (2014)
We summarize the identification proof from Kuroki and Pearl (2014), relying on Lemma B.1.
Lemma B.1.
$\mathbf{M}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{-1},$
Proof.
[…]
$\mathbf{P}_{y,W\mid Z,a}=\mathbf{P}_{W\mid U}\operatorname{diag}(\mathbf{P}_{y\mid U,a})\mathbf{P}_{U\mid Z,a}.$ (9)
By Assumption 10, $\mathbf{P}_{U\mid Z,a}^{-1}$ exists and
$\mathbf{P}_{W\mid Z,a}^{-1}=\mathbf{P}_{U\mid Z,a}^{-1}\,\mathbf{P}_{W\mid U}^{-1}.$
Hence,
$\mathbf{P}_{y,W\mid Z,a}\mathbf{P}_{W\mid Z,a}^{-1}=\mathbf{P}_{W\mid U}\operatorname{diag}(\mathbf{P}_{y\mid U,a})\mathbf{P}_{W\mid U}^{-1}.$ (10)
Then Lemma B.1 allows us to identify the column vectors of $\mathbf{P}_{W\mid U}$ under Assumption 11. In other words, we recover conditional distributions $f_{W\mid U}(w\mid u_{i})$, up to the labeling of the latent variable $U$. Inverting $f_{Y,W\mid A}(y,w\mid a)=\sum_{i}f_{Y,U\mid A}(y,u_{i}\mid a)\,f_{W\mid U}(w\mid u_{i}),$ we recover
$f_{Y\mid U,A}(y\mid u_{i},a),\quad f_{U}(u_{i}\mid a).$
Further by $W\perp U\mid A$, latent factors for label $i$ recovered across values of $a\in\mathcal{A}$ can be aligned, giving the components comprising the adjustment formula in Equation (1). ∎
Appendix C Mediator Proxy Analogue for Kuroki and Pearl (2018)
@@
$f_{Y(a)}(y)$ is identified by the proof in Appendix B, replacing $U$ with $M$.